# Topological excitations in a spin-orbit coupled spin-1 Bose-Einstein condensate under sinusoidally varying magnetic fields

### v1

@@ Abstract
We present a theoretical study of a spin-orbit coupled spin-1 Bose-Einstein condensate under the influence of sinusoidally varying magnetic fields. In the ground state of the ferromagnetic spin-1 condensate, we investigate topological excitations in the system arising due to the combined effect of Rashba spin-orbit coupling and an in-plane sinusoidally varying magnetic field. In this work, we offer a comparative study for various choices of magnetic fields in the $x\rm{-}y$ plane. For a fixed field strength, the spin-orbit coupled system sustains a rich variety of exotic vortex structures ranging from vortex-antivortex lattices to vortex clusters as we increase the coupling strength.
1 Introduction
@@
The experimental realization of spin-orbit coupling (SOC) in ultracold atomic physics [1, 2, 3, 4, 5, 6, 7, 8, 9] stands as a transformative milestone, opening new avenues for exploration at the quantum frontier. This groundbreaking achievement propels the field into uncharted territories where the delicate interplay between particle spins and their orbital motion can be precisely controlled and manipulated [10, 11, 12]. Typically, the order parameter of the spin-orbit (SO) coupled trapped Bose-Einstein condensates (BECs) [3] exhibit a very rich variety of exotic quantum states such as vortex lattice [13, 14], half-quantum vortex [3, 15], vortex neclace [16], skyrmion [17, 18, 19], domain wall [2], solitons [20, 21, 22], knots [23, 24], supersolid stripe phase [25, 26, 27], spin and field squeezing [28, 29] etc. Moreover in recent times, research has been extended in new directions such as SO-induced supersolidity [30, 27], spin-tensor momentum coupling [31], spin-orbital angular momentum coupling [32], etc. The introduction of such couplings has led to many interesting quantum states with novel topological textures [33, 34]. Recently, the splitting of vortex cores [35] and the first-order phase transitions [36] have been experimentally observed. These topological defects in ultracold condensates are vital for drawing analogies with those of condensed matter systems.
@@
In this light, spinor condensates [37, 38, 11, 39] emerge as exceptional platforms within ultracold atomic physics, serving as a prolific area of research for the investigation of solid-state-like properties [40]. These remarkable quantum systems, exhibiting a multitude of spin degrees of freedom, provide an intriguing bridge between the microscopic world of ultracold atoms and the condensed matter physics of solids. Spinor condensates offer a controlled environment to explore phenomena such as magnetic ordering [41], topological excitations [21] and defects [42], and quantum phase transitions [43, 44], shedding light on the intricate connection between quantum gases and condensed matter systems. This convergence of ultracold atomic physics and solid-state physics not only enriches our understanding of fundamental physics but also holds the promise of unveiling new quantum phenomena with potential applications in quantum information processing and beyond [45].
@@
On the other hand, recent literature shows that in-plane magnetic fields [46, 47, 48, 49, 50, 51, 52, 53, 54, 55] can be significantly used in investigating the equisite many-body physics of ultracold atoms. Relevant studies demonstrate that in-plane gradient magnetic fields can be used to achieve SOC and spin Hall states in BECs [56]. These magnetic fields play a vital role in creating synthetic non-Abelian gauge fields [57, 58]. Varying magnetic fields can be accredited for the observation of a wide assortment of topological defects, such as magnetic monopoles [50, 59], quantum knots [23], skyrmions [18], etc. The ground state of spinor BECs in the gradient field has been investigated to exhibit the central Mermin-Ho vortex with symmetrical vortex lattices [60]. The combined effect of SOC and gradient magnetic field on such systems resulted in skyrmion chains [18, 61] in the ground state order parameter. To the best of our knowledge, the effect of a sinusoidally varying magnetic field on SO-coupled BECs has not been given due attention. In this work, we study the combined effect of a sinusoidally varying magnetic field and SOC on the ground state of a spin-1 BEC. Here, we consider an SO-coupled spin-1 BEC subject to various combinations of sinusoidally varying magnetic fields in th $x\rm{-}y$ plane. We offer a comparative study between the various choices of sinusoidally varying magnetic fields, thereby highlighting the different topological structures so obtained for each case.
@@
This paper is organised as follows: In section 2, we introduce our model system and discuss its mean-field ground state under different situations for the different combinations of sinusoidally varying in-plane magnetic field. In section 3, we discuss the results obtained by numerically solving the Gross-Pitaevskii (GP) equations and investigate the topological structures (vortices) generated in the system as a consequence of the interplay between SOC, sinusoidally varying magnetic field with the spin-1 order parameter. In section 4, we summarize our findings and provide an outreach for future scope.
2 Model
@@
In the mean-field, we consider a quasi-2D Rashba SO-coupled ferromagnetic spin-1 BEC of $N$ identical atoms in an harmonic trap in a sinusoidally varying magnetic field. The effective Hamiltonian [62] in the Gross-Pitaevskii form can be given as $H=H_{0}+H_{int}$,
@@
The Rashba SOC interaction [1, 3] is given as $\nu_{SOC}=\gamma(f_{x}p_{y}-f_{y}p_{x})$ where $\gamma$ is the SOC strength and $(p_{x},p_{y})$ represents the momentum in the quasi-2D space. Note that this SOC is SU(2) type, therefore the states $\Psi_{1}$ and $\Psi_{-1}$ cannot be coupled directly. In real experiments, the SU(2) SOC Hamiltonian can be achieved by Raman coupling three laser beams of different polarizations and frequencies. The external magnetic field is given as $B(\mathbf{r})=(B_{x}\hat{x}-B_{y}\hat{y})$ in the 2D plane. Here, $B_{x},B_{y}$ are sinusoidally varying magnetic fields along the $x$ and $y$ directions. We absorb the negative sign in the direction and denote $B(\mathbf{r})=(B_{x},B_{y})$. The Lande’s g-factor $g_{F}=-\frac{1}{2}$ and $\mu_{B}$ is the Bohr’s magneton.
@@
$\rm{i}\frac{\partial{\psi_{1}}}{\partial{t}}$ $=\left[-\frac{1}{2}\nabla^{2}+V+c_{0}|\psi|^{2}+c_{2}\left(|\psi_{1}|^{2}+|\psi_{0}|^{2}-|\psi_{\rm{-}1}|^{2}\right)\right]\psi_{1}+c_{2}\psi_{\rm{-}1}^{*}\psi_{0}^{2}$
$+(B_{x}+\rm{i}B_{y})\psi_{0}-\rm{i}{\gamma}\left(\frac{\partial}{\partial y}+\rm{i}\frac{\partial}{\partial x}\right)\psi_{0}$ (3)
$\rm{i}\frac{\partial{\psi_{0}}}{\partial{t}}$ $=\left[-\frac{1}{2}\nabla^{2}+V+c_{0}|\psi|^{2}+c_{2}\left(|\psi_{1}|^{2}+|\psi_{\rm{-}1}|^{2}\right)\right]\psi_{0}+2c_{2}\psi_{1}\psi_{0}^{*}\psi_{\rm{-}1}$
$+(B_{x}-\rm{i}B_{y})\psi_{1}+(B_{x}+\rm{i}B_{y})\psi_{\rm{-}1}-\rm{i}{\gamma}\left(\frac{\partial}{\partial y}-\rm{i}\frac{\partial}{\partial x}\right)\psi_{1}-\rm{i}{\gamma}\left(\frac{\partial}{\partial y}+\rm{i}\frac{\partial}{\partial x}\right)\psi_{\rm{-}1}$ (4)
$\rm{i}\frac{\partial{\psi_{\rm{-}1}}}{\partial{t}}$ $=\left[-\frac{1}{2}\nabla^{2}+V+c_{0}|\psi|^{2}+c_{2}\left(-|\psi_{1}|^{2}+|\psi_{0}|^{2}+|\psi_{\rm{-}1}|^{2}\right)\right]\psi_{\rm{-}1}+c_{2}\psi_{1}^{*}\psi_{0}^{2}$
$+(B_{x}-\rm{i}B_{y})\psi_{0}-\rm{i}{\gamma}\left(\frac{\partial}{\partial y}-\rm{i}\frac{\partial}{\partial x}\right)\psi_{0}$ (5)
where $\psi_{m}=N^{-1/2}l\Psi_{m}\quad(m=\pm 1,0)$ denotes the dimensionless $m$th component wave function and $|\psi|^{2}=\sum_{m}|\psi_{m}|^{2}$ is the total particle density confined in dimensionless external potential $V=(x^{2}+y^{2})/2$. The contact interaction parameters in the dimensionless form are given as $c_{0}=2N\sqrt{2\pi\lambda}(a_{0}+2a_{2})/3l$ and $c_{2}=2N\sqrt{2\pi\lambda}(a_{2}-a_{0})/3l$, respectively. $B,{\gamma}$ are the dimensionless magnetic field and SOC strengths, respectively. In our numerical calculations, the lengths, time, energy(interaction and SOC) and magnetic field are measured in the units of $l,1/\omega_{\perp},\hbar\omega_{\perp}$ and $(g_{F}\mu_{B}l)/\hbar\omega_{\perp}$, respectively.
The vorticity [63] of the $m-$th component can be measured as
$\Omega_{m}=\nabla\times J_{m}$ (6)
@@
$J_{m}=\frac{\rm{i}\hbar}{2M}(\psi_{m}\nabla\psi^{*}_{m}-\psi^{*}_{m}\nabla\psi_{m})$ (7)
The topological excitation of the sytem arising due to the formation of vortices is characterised by the spin texture [64]. For a spin-1 BEC, the spin texture can be written as [65, 66],
$S_{\alpha}=\sum_{m,n=0,\pm 1}\psi^{*}_{m}(f_{\alpha})_{m,n}\psi_{n}/|\psi|^{2}\quad(\alpha=x,y,z)$ (8)
The topological charge density which characterises the topological structure of the system is given as,
$q(\mathbf{r})=\frac{1}{4\pi}\mathbf{s}\cdot\left(\frac{\partial\mathbf{s}}{\partial x}\times\frac{\partial\mathbf{s}}{\partial y}\right)$ (9)
and the total topological charge $Q$ is defined as
@@
Here $\mathbf{s}=S/|S|$ and the topological charge $|Q|$ is unchanged under any exchange of $S_{x},S_{y},S_{z}$.
3 Results and Discussions
@@
We investigate the ground state structures of the spin-orbit coupled ferromagnetic spin-1 BEC under the influence of external in-plane magnetic field. In this section, we first discuss the effect of a polarized magnetic field in the $x$ or $y$ direction. The ground state density profile and the spin textures are shown for cases both with and without SOC. Thereafter we discuss cases for sinusoidally varying in-plane magnetic fields. For each combination of external magnetic field $B(\mathbf{r})=(B_{x},B_{y})$, we present the ground state structures with and without SOC.
@@
We offer a detailed numerical study to analyze the effects of system parameters on the ground state structures. The luxury of the present system lies in the large number of tunable parameters, which includes the spin-independent and spin-exchange contact interactions, SOC strength, magnetic field strength. To focus our attention on the effects of SOC and in-plane magnetic field, we have considered $10^{4}$ number of $\rm{{}^{87}{Rb}}$ atoms confined in a quasi-2D harmonic trap $V(\mathbf{r})$ with transverse trapping frequency, $\omega_{\perp}=2\pi\times 20$ Hz with trap aspect ratio, $\lambda=\omega_{z}/\omega_{\perp}=20$. The $s\rm{-}$wave scattering lengths for channels of total spin 0 and 2 are $a_{0}=101.8a_{B}$ and $a_{2}=100.4a_{B}$ respectively, where $a_{B}$ is the Bohr radius. With these parameters [67], using the split-step Crank-Nicolson scheme [68, 69, 70], we numerically solve the governing GP equations (2-2) for imaginary time propagation to generate the ground state solutions for diverse choices of magnetic fields for numerous SO coupling strengths. All of our simulations run from a spatial extent of $-20l$ to $20l$ in both $x$ and $y$ directions with $2001\times 2001$ grid points. The employed spatial discretization refers to $\Delta x=\Delta y=0.02l$, with time step of numerical integration $\delta t=(2\times 10^{-4})/\omega_{\perp}$.
3.1 Polarized gradient magnetic field
[…]

### v2

# Sinusoidal magnetic field induced topological excitations in a spin-orbit coupled spinor condensate

@@ Abstract
We explore topological excitations in a spin-1 Bose-Einstein condensate subjected to an in-plane sinusoidally varying magnetic field and Rashba spin-orbit coupling (SOC). In the absence of SOC, the periodic magnetic field induces vortex-anti-vortex structures in the |$\ket{F=1,m_{F}=\pm 1}$ condensates at saddle-points, such that the net topological charge remains zero. The introduction of Rashba SOC breaks the system’s symmetry, leading to non-conservation of overall angular momentum in the spin-1 condensate. This anisotropy results in the emergence of certain skyrmion spin textures. We provide a comparative study for various in-plane magnetic field configurations while keeping the SOC strength constant. Our numerical simulations within the mean-field framework reveal the potential to engineer diverse topological excitations controlled by the interplay between spin-orbit coupling and in-plane magnetic field in a spinor condensate.
I Introduction
@@
The experimental achievement of introducing spin-orbit coupling (SOC) in ultracold atomic physics [1, 2, 3, 4, 5, 6, 7, 8, 9] marks a transformative milestone, opening up new frontiers for exploration in the quantum realm. This groundbreaking progress propels the field into unexplored territories, allowing for precise control and manipulation of the intricate interplay between particle spins and orbital motion [10, 11, 12, 13, 14, 15]. Spin-orbit (SO) coupled trapped Bose-Einstein condensates (BECs) typically exhibit a diverse range of exotic quantum states in their order parameter, including vortex lattice [16, 17, 18], half-quantum vortex [3, 19, 20], vortex neclace [21], skyrmion [22, 23, 24, 25], domain wall [2], solitons [26, 27, 28, 29, 30], knots [31, 32], supersolid stripe phase [33, 34, 35, 36, 37], SO-induced supersolidity [38, 37], spin and field squeezing [39, 40], etc. Furthermore, recent research has extended into novel directions such as spin-tensor momentum coupling [41, 42, 43, 44, 45, 46, 47], spin-orbital angular momentum coupling [48, 49, 50, 51, 52, 53, 54, 55, 56, 57], and more. Notably, there have been recent experimental observations of the splitting of vortex cores [58] and the first-order phase separations [34, 59, 60, 61]. These topological defects in ultracold condensates play a crucial role in drawing analogies with those found in condensed matter systems.
@@
In this context, spinor condensates [62, 63, 64, 65, 66, 67] emerge as exceptional platforms within ultracold atomic physics, serving as a copious area of research for the investigation of solid-state-like properties [68, 69]. These remarkable quantum systems, exhibiting a multitude of spin degrees of freedom, provide an intriguing bridge between the microscopic world of ultracold atoms and the condensed matter physics of solids. Spinor condensates offer a controlled environment to explore phenomena such as magnetic ordering [70], topological excitations and defects [71, 29] turbulence [72, 73, 74], shedding light on the intricate connection between quantum gases and condensed matter systems [71]. This synergy of ultracold atomic physics and solid-state physics not only enriches our understanding of fundamental physics but also holds the promise of unveiling new quantum phenomena with potential applications in quantum information processing and beyond [75].
@@
On the other hand, recent literature shows that in-plane magnetic fields [76, 77, 78, 79, 80, 81, 82, 83, 84, 85] can be significantly used in investigating the equisite many-body physics of ultracold atoms. Relevant studies demonstrate that in-plane gradient magnetic fields can be used to achieve SOC and spin Hall states in BECs [86]. These magnetic fields play a vital role in creating synthetic non-Abelian gauge fields [87, 88]. Varying magnetic fields can be accredited for the observation of a wide assortment of topological defects, such as magnetic monopoles [80, 89], quantum knots [31], skyrmions [24]. The ground state of spinor BECs in the gradient field has been investigated to exhibit the central Mermin-Ho vortex with symmetrical vortex lattices [90, 91]. The combined effect of SOC and gradient magnetic field on such systems result in skyrmion chains [24, 92] in the ground state.
Previous investigations on SO-coupled spinor BECs have primarily focused on the effects of either a uniform magnetic field or a polarized-gradient magnetic field. However, the influence of other spatially varying magnetic fields have not been addressed adequately. The inclusion of sinusoidal magnetic fields, although less common than uniform or gradient fields, the experiment [93] provides a unique protocol to induce spatially varying magnetic interactions. These fields can create periodic modulations in the spin-dependent interactions, leading to the emergence of novel quantum states and enhanced control over spin textures.
In this work, we investigate the ground state of a spin-1 BEC subjected to diverse combinations of sinusoidally varying magnetic fields within the $x$-$y$ plane so as to explore the synergistic impact of a periodic magnetic field and the SOC. The intricate interplay between the external magnetic field and SOC yields topological exciations in the spin-1 BEC in the form of exotic vortices and spin textures. We offer a comparative study between the various choices of sinusoidally varying magnetic fields, thereby highlighting the different topological structures so obtained for each case. Our study aims to leverage the sinusoidal magnetic fields to explore and stabilize complex spin textures such as skyrmions, which are topologically protected spin configurations characterized by a whirling pattern of spins. Skyrmions are particularly compelling due to their robustness against perturbations and their potential as information carriers for next-generation spintronic memory and logic computing devices with high-density, offering non-volatile storage, low power consumption [94, 95].
@@
This paper is organised as follows: In section II, we introduce our model system and discuss its mean-field ground state under different situations for the different combinations of sinusoidally varying in-plane magnetic field, and thereby, casting these into the Gross-Pitaevskii (GP) equations. In section III, we discuss the results obtained by numerically solving the GP equations and investigate the topological structures (vortices and spin textures) generated in the system as a consequence of the interplay between the SOC and the oscillating magnetic field with the spin-1 order parameter. In section IV, we summarize our findings and provide insights for future scope. We discuss the impact of varying the SOC strength in appendix A.
II Model
@@
In the mean-field, we consider a Rashba SO-coupled ferromagnetic spin-1 BEC of $N$ atoms in a quasi-2D harmonic trap in a spatially varying magnetic field. The effective Hamiltonian [96] in the Gross-Pitaevskii form can be given as $H=H_{0}+H_{\rm int}$,
@@
The Rashba SOC interaction [1, 3] is given as $\nu_{\rm SOC}=\gamma(f_{x}p_{y}-f_{y}p_{x})$, where $\gamma$ is the SOC strength and $(p_{x},p_{y})$ represents the momentum in the quasi-2D space. Note that this SOC is SU(2) type, therefore it cannot couple thr states $\Psi_{1}$ and $\Psi_{-1}$, directly. In real experiments, the SU(2) SOC Hamiltonian can be achieved by Raman coupling three laser beams of different polarizations and frequencies [97]. In addition, the external magnetic field is given as $B(\mathbf{r})={(B_{x}\hat{x}-B_{y}\hat{y})}$ in the 2D plane. Here, $B_{x},B_{y}$ are spatially varying magnetic fields along the $x$ and $y$ directions. We absorb the negative sign in the direction and denote $B(\mathbf{r})={(B_{x},B_{y})}$. The Lande’s g-factor $g_{F}=-\frac{1}{2}$ and $\mu_{B}$ is the Bohr magneton.
@@
${i}\frac{\partial{\psi_{1}}}{\partial{t}}$ $=\bigg{[}-\frac{1}{2}\nabla^{2}+V+c_{0}|\psi|^{2}+c_{2}\bigg{(}|\psi_{1}|^{2}+|\psi_{0}|^{2}$
$-|\psi_{{-}1}|^{2}\bigg{)}\bigg{]}\psi_{1}+c_{2}\psi_{{-}1}^{*}\psi_{0}^{2}+({B_{x}}+{i}B_{y})\psi_{0}$
$-{i}{\gamma}\left(\frac{\partial}{\partial y}+{i}\frac{\partial}{\partial x}\right)\psi_{0}$ (3)
$i\frac{\partial{\psi_{0}}}{\partial{t}}$ $=\left[-\frac{1}{2}\nabla^{2}+V+c_{0}|\psi|^{2}+c_{2}\left(|\psi_{1}|^{2}+|\psi_{{-}1}|^{2}\right)\right]\psi_{0}$
$+2c_{2}\psi_{1}\psi_{0}^{*}\psi_{{-}1}+(B_{x}-iB_{y})\psi_{1}+(B_{x}+iB_{y})\psi_{{-}1}$
$-i{\gamma}\left(\frac{\partial}{\partial y}-i\frac{\partial}{\partial x}\right)\psi_{1}-i{\gamma}\left(\frac{\partial}{\partial y}+i\frac{\partial}{\partial x}\right)\psi_{{-}1}$ (4)
$i\frac{\partial{\psi_{\rm{-}1}}}{\partial{t}}$ $=\bigg{[}-\frac{1}{2}\nabla^{2}+V+c_{0}|\psi|^{2}+c_{2}\bigg{(}-|\psi_{1}|^{2}+|\psi_{0}|^{2}$
$+|\psi_{{-}1}|^{2}\bigg{)}\bigg{]}\psi_{{-}1}+c_{2}\psi_{1}^{*}\psi_{0}^{2}+({B_{x}}-iB_{y})\psi_{0}$
$-i{\gamma}\left(\frac{\partial}{\partial y}-i\frac{\partial}{\partial x}\right)\psi_{0}$ (5)
where $\psi_{m}=N^{-1/2}l\Psi_{m}\quad(m=\pm 1,0)$ denotes the dimensionless $m$-th component wave function and $|\psi|^{2}=\sum_{m}|\psi_{m}|^{2}$ is the total particle density confined in dimensionless external potential $V=(x^{2}+y^{2})/2$. The contact interaction parameters in the dimensionless form are given as $c_{0}=2N\sqrt{2\pi\lambda}(a_{0}+2a_{2})/3l$ and $c_{2}=2N\sqrt{2\pi\lambda}(a_{2}-a_{0})/3l$, respectively. In our numerical calculations, the lengths, time, energy (interaction and SOC) and magnetic field are measured in the units of $l,1/\omega_{\perp},\hbar\omega_{\perp}$ and $(g_{F}\mu_{B}l)/\hbar\omega_{\perp}$, respectively.
The vorticity of the $m$-th component can be measured as [98, 99]
$\Omega_{m}=\nabla\times J_{m},\quad m=\pm 1,0$ (6)
@@
$J_{m}=\frac{{i}\hbar}{2M}(\psi_{m}\nabla\psi^{*}_{m}-\psi^{*}_{m}\nabla\psi_{m}).$ (7)
The topological excitation of the system arising resulting in the formation of vortices, is characterised by the spin texture [100]. In a spinor BEC, the spin texture describes the spatial variation of the spin degrees of freedom across the condensate, ranging from complex configurations like skyrmions and spin vortices. Grasping these spin textures is key to delving into the intricate physics of spinor BECs. For a spin-1 BEC, the spin texture can be written as [101, 102],
$\mathbf{S_{\alpha}}=\sum_{m,n=0,\pm 1}\psi^{*}_{m}(f_{\alpha})_{m,n}\psi_{n}/|\psi|^{2}\quad(\alpha=x,y,z).$ (8)
The topological charge density which characterises the topological spin structure of the system is given as,
$q(\mathbf{r})=\frac{1}{4\pi}\mathbf{s}\cdot\left(\frac{\partial\mathbf{s}}{\partial x}\times\frac{\partial\mathbf{s}}{\partial y}\right).$ (9)
Here $\mathbf{s}=\mathbf{S/|S|}$ and the total topological charge $Q$ is defined as
@@
The topological charge $|Q|$ is unchanged, however one exchanges the components ${\mathbf{S}_{x},\mathbf{S}_{y},\mathbf{S}_{z}}$.
III Results and Discussions
@@
We investigate the ground state structures of the spin-orbit coupled ferromagnetic spin-1 BEC under the influence of external in-plane magnetic field. In this section, we first discuss the effect of a polarized gradient magnetic field in the $x$ or $y$ direction. We show the ground state density profiles and the spin textures for cases both with and without SOC. Thereafter, we discuss cases for sinusoidally varying in-plane magnetic fields. Further, we present the ground state structures with and without SOC for each combination of external magnetic field $B(\mathbf{r})={(B_{x},B_{y})}={B_{0}(f(x),g(y))}$, where ${B_{0}}$ is the strength of external magnetic field in dimensionless units.
@@
We offer a detailed numerical study to analyze how the system parameters affect the ground state structures. The richness of the present system lies in the large number of tunable parameters, which includes the spin-independent and spin-exchange contact interactions $c_{0}$ and $c_{2}$ respectively, SOC strength, $\gamma$ and magnetic field strength, ${B_{0}}$. Here, we have considered $10^{4}$ number of $\rm{{}^{87}{Rb}}$ atoms confined in a quasi-2D harmonic trap $V(\mathbf{r})$ with transverse trapping frequency, $\omega_{\perp}=2\pi\times 20$ Hz with trap aspect ratio, $\lambda=\omega_{z}/\omega_{\perp}=20$ [103] and the characteristic oscillator length, $l=2.41\mu m$. The $s$-wave scattering lengths for channels of total spin 0 and 2 are $a_{0}=101.8a_{B}$ and $a_{2}=100.4a_{B}$, respectively, where $a_{B}$ is the Bohr radius. We keep the magnetic field strength ${B_{0}}=0.25$ for all choices of magnetic fields. And for all the cases with SOC, we keep the SOC strength fixed at $\gamma=0.5$. With these parameters, using the split-step Crank-Nicolson scheme [104, 105, 106], we numerically solve the governing GP equations (II-II) for imaginary time propagation to generate the ground state solutions in the presence of various in-plane polarized and cross-polarized magnetic fields for a fixed SO coupling strength. All of our simulations run from a spatial extent of $-20l$ to $20l$ in both $x$ and $y$ directions with $2001\times 2001$ grid points. The employed spatial discretization refers to $\Delta x=\Delta y=0.02l$, with a time step $\delta t=(2\times 10^{-4})/\omega_{\perp}$.
III.1 In-plane polarized gradient magnetic field